# Spectral Flow Learning Theory: Finite-Sample Guarantees for Vector-Field Identification

### v3

@@ Abstract
We study the identification of continuous-time vector fields from irregularly sampled trajectories. We introduce Spectral Flow Learning (SFL), which learns in a windowed flow space using a lag-linear label operator that aggregates lagged Koopman actions. We provide finite-sample high-probability (FS-HP) guarantees for the class of variable-step linear multistep methods (vLLM). The FS-HP rates are constructed using spectral regularization with qualification-controlled filters for flow predictors under standard source and filter assumptions. A multistep observability inequality links flow error to vector-field error and yields two-term bounds that combine a statistical rate with an explicit discretization bias from vLMM theory.
This preliminary preprint states the results and sketches proofs, with full proofs and extensions deferred to a journal version.
@@ I Introduction
In this manuscript, we develop preliminary components of a Spectral Flow Learning (SFL) theory, a spectral-regularization view of learning dynamical laws from irregular trajectories. In SFL, learning is posed as an inverse problem [1] driven by the inclusion map $\mathsf{I}$ of a vector-valued Reproducing Kernel Hilbert Space (vvRKHS) into a population $L^{2}$ space; the covariance operator ${\mathsf{T}}={\mathsf{I}}^{*}{\mathsf{I}}$, a spectral filter ${\mathsf{g}}_{\lambda}$, its qualification $\nu$, and a Lipschitz exponent $\mu$ govern approximation/stability rates [2, 3].
We take the flow as the primary regression object and lift the finite-sample high-probability (FS-HP) excess risk bound from flow space to the vector field via an observability inequality induced by a Koopman-lag forcing operator $\mathsf{B}$. Our operator notation and bias-variance decomposition follow [4]. Deterministic modeling error from discretization is controlled by variable-step size linear multistep method (vLMM) theory, bounded coefficients, variable-step consistency, and ODE regularity condition [5, Ch. 3.5].
@@ II Setup and Problem Formulation
This section makes the setup precise. We first fix notation and basic functional-analytic conventions. We then specify the dynamical system and its flows (autonomous and input-driven), introduce the design law that governs state visitation and the step law that governs time advances on irregular grids. Finally, we state the finite-sample, high-probability bounds to be proved later (Eq. (1)). These choices determine the population spaces and bounded operators used throughout.
@@ II-A Notation
$\mathbb{R}$ denotes the real number line. Vectors in $\mathbb{R}^{n}$ are column vectors. For a set $A$, $\mathbf{1}_{A}$ denotes its indicator function. For a test function $\varphi$ and a measure $\mu$, the pairing is defined by $\langle\varphi,\mu\rangle\coloneqq\int\varphi d\mu.$ We write $a\lesssim b$ when there exists a universal constant $C>0$ with $a\leq Cb$, and we write $a\asymp b$ when $a\lesssim b$ and $b\lesssim a$. For a Hilbert space $\mathcal{H}$, $\langle u,v\rangle_{\mathcal{H}}$ denotes the inner product and $\|u\|_{\mathcal{H}}$ denotes the norm. On $\mathbb{R}^{n}$ the inner product is $\langle x,y\rangle=x^{\top}y$ and the Euclidean norm is $\|x\|_{2}=(x^{\top}x)^{1/2}$. The space $L^{2}(\mathcal{X},\mu)$ consists of square-integrable functions and its norm is $\|f\|_{L^{2}}=\sqrt{\int|f|^{2}d\mu}.$ The space $\mathcal{C}(\mathcal{X})$ denotes continuous real-valued functions on $\mathcal{X}$, and $\mathcal{C}^{k}(\mathcal{X})$ denotes the $k$-times continuously differentiable functions. The Banach space $\mathcal{L}(\mathcal{H}_{1},\mathcal{H}_{2})$ denotes bounded linear operators $\mathsf{T}:\mathcal{H}_{1}\to\mathcal{H}_{2}$, and the operator norm is $\|\mathsf{T}\|_{\mathcal{L}}=\sup_{\|x\|=1}\|\mathsf{T}x\|.$ The adjoint of $\mathsf{T}$ is written $\mathsf{T}^{*}$, and $\mathrm{Id}$ denotes the identity operator. An operator is self-adjoint when $\mathsf{T}=\mathsf{T}^{*}$, and it is positive semidefinite when $\mathsf{T}\succeq 0$. For a self-adjoint operator $\mathsf{T}$, $\lambda_{\min}(\mathsf{T})$ and $\lambda_{\max}(\mathsf{T})$ denote its minimal and maximal eigenvalues. The set $\mathcal{P}(\mathcal{X})$ denotes Borel probability measures on $\mathcal{X}$. For $\rho\in\mathcal{P}(\mathcal{X})$, $\operatorname{supp}(\rho)$ denotes the support and $\delta_{x}$ denotes the Dirac mass at $x$. We write $\mathbb{E}[\cdot]$ for expectation and $\mathbb{E}[\ \cdot\mid\cdot\ ]$ for conditional expectation. We write $a_{n}=\mathcal{O}(b_{n})$ when $\sup_{n}|a_{n}/b_{n}|<\infty$.
@@ II-B Dynamical System and Flows
In SFL, the flow is the primary regression object: we view the problem of learning vector-fields from trajectory data as identifying the one–step map $(x,h)\mapsto\Phi^{h}(x)$ in an $L^{2}$ space and then lift the flow estimator to the vector field. This choice makes the feature-label pairs well-posed under irregular sampling and enables us to apply spectral regularization methods in learning theory to system identification problems. With this motivation, we now define the dynamical model and its flows. Let $\mathcal{X}\subset\mathbb{R}^{n}$ be compact and $\mathcal{B}=\mathcal{B}(\mathcal{X})$ its Borel $\sigma$-algebra. Consider the control‑affine dynamics:
@@
where $\tau^{s}u$ is the time-shift operator of the input signal $u$ by $s$, such that $(\tau^{s}u)(t)\coloneqq u(t+s),\forall t\geq 0$. We write $\Phi_{u}^{t}\coloneqq\Phi_{u}^{t,0}$. Special case: $u\equiv 0\Rightarrow\Phi_{0}^{t}=\phi^{t}$. We denote flows generated by an autonomous vector field $f$ and control-affine dynamics $(f,g)$ as $\phi^{h}(\ \cdot\ ;f)$ and $\Phi^{h}_{u}(\ \cdot\ ;f,g)$ respectively. Furthermore, we use $\Phi^{h}(\cdot)$ when the discussion is agnostic with respect to the existence of a system input $u$, and $\Phi^{h}_{\rho}(\cdot)$ when the input is determined by a design law $\rho$.
@@ II-C Design Law (state visitation distribution)
The design law $\rho_{\mathcal{X}}^{\mathrm{design}}\in\mathcal{P}(\mathcal{X})$, specifies where the experiment spends time in $\mathcal{X}$. $\rho_{\mathcal{X}}^{\mathrm{design}}$ is a Borel probability measure on $(\mathcal{X},\mathcal{B})$. There are multiple ways to construct a design law; we are interested in the continuous-time time-average law. Along a single path $x(t)=\Phi_{u}^{t}(x_{0})$:
@@
where $\delta_{x}$ is the Dirac measure: $\delta_{x}(A)=\mathbf{1}_{A}(x)$, $A\in\mathcal{B}$. The design law $\rho_{\mathcal{X}}^{\mathrm{design}}$ determines coverage of states; identifiability over the identificaiton domain $\mathcal{X}_{\rm ID}$ requires $\mathcal{X}_{\rm ID}\subseteq\operatorname{supp}(\rho_{\mathcal{X}}^{\mathrm{design}})$.
@@
The design law induces the geometry in which we measure errors and therefore governs identifiability. In particular, we can view identifiability as an input–design problem: choose an input policy $u(\cdot)$ so that the induced $\rho_{\mathcal{X}}^{\mathrm{design}}(u)$ places mass on the domain of interest. Moreover, the population risks $\mathcal{E}(\Phi_{\rho})$ and $\mathcal{E}(f_{\rho})$ used later in the problem formulation are squared $L^{2}$-errors with respect to the design weights $\rho^{\rm design}_{\mathcal{X}}\times\rho_{\mathbf{h}}$ and $\rho^{\rm design}_{\mathcal{X}}$. The requirement of $\mathcal{X}_{\rm ID}\subseteq\operatorname{supp}(\rho_{\mathcal{X}}^{\mathrm{design}})$ ensures that discrepancies on the identification domain are penalized and thus identifiable.
@@ II-D Step Law (time‑advance distribution)
Let $\boldsymbol{h}\coloneqq(0,h_{\max}]$, the step law $\rho_{\boldsymbol{h}}\in\mathcal{P}\big((0,h_{\max}]\big)$ specifies how far we advance in time when forming one‑step labels $(x_{k},h_{k})\mapsto x_{k+1}$. It is a probability measure on the interval $(0,h_{\max}]$. In this work, we are interested in a deterministic schedule $\{h_{k}\}$. Specifically, a step law $\rho_{\boldsymbol{h}}$ can be constructed through index‑average:
$\lim_{N\to\infty}\bigg\langle\varphi,\frac{1}{N}\sum_{k=0}^{N-1}\delta_{h_{k}}\bigg\rangle=\langle\varphi,\rho_{\boldsymbol{h}}\rangle,\quad\forall\ \varphi\in\mathcal{C}(\mathcal{X}).$
We are ready to formulate the problem of identifying flow and vector field FS-HP bounds in the SFL framework.
@@ II-E Problem Formulation (FS-HP bound)
$f_{\rho}(x)\coloneqq\lim_{r\downarrow 0}\;\mathbb{E}\!\left[\frac{Y-X}{H}\bigg|X=x,0<H\leq r\right],$
Under the flow source condition $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$, [4, Eq. 9], and a spectral filter ${\mathsf{g}}_{\lambda}$ with qualification $\nu\geq r$ and Lipschitz exponent $\mu$ (set $\beta\coloneqq\max\{1,2\mu\}$), [4, Def. 1], our goal is to establish finite-sample, high-probability (FS–HP) bounds of the form:
@@
$\underbrace{\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}}_{\text{vector-field excess risk}}$ $\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}}\log\!\frac{4}{\eta}\ell^{-\frac{2r}{2r+\beta}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}}\mathbb{E}[H^{2p+2}],$ (2)
which holds with probability at least $1-\eta$ over the model $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}$. Fig. 1 shows the road-map to proof the FS-HP bounds for the flow excess risk in Theorem 1 and the FS-HP bounds for the vector field excess risk in Theorem 2.
vLMM Assumptions Spectral Regularization Asm. III.1 bounded step ratios/coeffs ($\zeta,\alpha,\beta$) Filter qual./Lipschitz (Def. [4, Def. 1]) Asm. III.2 order-$p$ consistency Source cond. on vector field $\Phi_{\rho}\in\Omega^{(\mathcal{V})}_{r,R}$ Asm. III.3 ODE regularity Source cond. on flow $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ Lemma 2 Uniform LTE Lemma 1 Observability via $\mathsf{B}$ Thm. 1 Flow FS–HP bound Thm. 2 Vector-field FS–HP bound
Figure 1: Dependency graph linking assumptions, lemmas, theorems, corollaries, and propositions used in the SFL analysis.
@@ III Variable Step Linear Multistep Method (vLMM) Theory
This section records some vLMM facts we use to turn flow-space error into vector-field error within SFL. The role of vLMM in the SFL framework is twofold: (i) it supplies the windowed surrogate that links vector fields to flows through the multistep forcing/aggregation operators (enabling observability), and (ii) it controls the deterministic discretization bias via order-$p$ truncation guarantees. Since, for variable-step methods, consistency depends only on step-ratio and not on absolute time shifts, a step-window is often parameterized by the step-ratio vector $\zeta$ and an anchor step $h$ in the vLMM framework.

### v4

# Spectral Flow Learning Theory: Finite-Sample Guarantees for System Identification

@@ Abstract
We study the identification of continuous-time vector fields from irregularly sampled trajectories. We introduce spectral flow learning, which learns in a windowed flow space using a lag-linear label operator that aggregates lagged Koopman actions. We provide finite-sample, high-probability (FS-HP) guarantees for the class of variable-step linear multistep methods (vLMM). The FS-HP rates are constructed using spectral regularization with qualification-controlled filters for flow predictors under standard source and filter assumptions. A multistep observability inequality links flow error to vector-field error and yields two-term bounds that combine a statistical rate with an explicit discretization bias from vLMM theory. Simulations on a controlled mass-spring system corroborate the theory and clarify conditioning, step-sample tradeoffs, and practical implications.
@@ I Introduction
Classical continuous-time identification lacks operator-theoretic flow-field bounds [7], which motivates the proposed spectral flow learning (SFL) framework that offers FS-HP guarantees for identification pipelines for linear and nonlinear systems. SFL offers a spectral-regularization view of learning dynamical laws from irregular trajectories. In particular, learning is posed as an inverse problem [6], driven by the inclusion map $\mathsf{I}$ of a vector-valued Reproducing Kernel Hilbert Space (vvRKHS) into a population $L^{2}$ space. We provide operator-theoretic FS-HP guarantees for two broad classes of system-identification pipelines—(i) flow learning and (ii) vector-field learning.
Flow learning treats the one-step forward transition map $\Phi$ as the regression target, whereas vector-field learning targets the instantaneous dynamics $f$. We establish FS-HP generalization guarantees for one-step transition learners under irregular sampling, and then transfer these guarantees to vector-field estimators by (i) coupling $\Phi$ to $f$ through a backward-looking multistep window and (ii) disentangling a data-driven statistical rate from an explicit discretization bias induced by the step schedule. This two-layer analysis covers a broad spectrum of methods: flow-learning pipelines such as ARX, subspace identification, the DMD family, and GP-SSMs. The vector-field pipelines covered include GP vector-field regression.
Section II fixes notation, specifies the control-affine dynamics and their flows, introduces the design and step laws, and states the FS-HP targets that determine the population spaces and bounded operators. Section III records the ratio-parameterized multistep form, the bounded-coefficient/step-ratio and order-$p$ consistency assumptions, and proves a uniform local-truncation estimate that controls deterministic bias [8, Ch. 3.5]. Section IV assembles the vvRKHS hypotheses, inclusion/covariance/integral operators, and the spectral filter with qualification and Lipschitz properties, together with flow- and vector-field-level source conditions [5, 2]. Section V defines the label map and vLMM residual via Koopman-lag forcing and state aggregation, establishes an observability inequality, and combines these ingredients with spectral regularization to obtain excess-risk bounds in flow space and their lifting to vector field, cleanly separating statistical rates from discretization bias. Section VI evaluates representative pipelines on a controlled mass-spring system, illustrating the issue of benchmarking across source conditions and the fast-sampling paradox. Our operator notation and bias-variance decomposition follow [11].
@@ II Setup and Problem Formulation
This section makes the problem setup precise. We first define notation and basic functional-analytic conventions. We then specify the dynamical system and its flows (autonomous and input-driven), introduce the design and step law that governs state visitation time advances on irregular grids, respectively. Finally, we state the finite-sample, high-probability bounds to be proved in Eq. (1). These choices determine the population spaces and bounded operators used throughout.
@@ II-A Notation
$\mathbb{R}$ denotes the real number line. Vectors in $\mathbb{R}^{n}$ are column vectors. For a set $A$, $\mathbf{1}_{A}$ denotes its indicator function. For a test function $\varphi$ and a measure $\mu$, the pairing is defined by $\langle\varphi,\mu\rangle\coloneqq\int\varphi d\mu.$ We write $a\lesssim b$ when there exists a universal constant $C>0$ with $a\leq Cb$, and we write $a\asymp b$ when $a\lesssim b$ and $b\lesssim a$. For a Hilbert space $\mathcal{H}$, $\langle u,v\rangle_{\mathcal{H}}$ denotes the inner product and $\|u\|_{\mathcal{H}}$ denotes the norm. On $\mathbb{R}^{n}$, the inner product is $\langle x,y\rangle=x^{\top}y$ and the Euclidean norm is $\|x\|_{2}=(x^{\top}x)^{1/2}$. The space $L^{2}(\mathcal{X},\mu)$ consists of square-integrable functions and its norm is $\|f\|_{L^{2}}=\sqrt{\int|f|^{2}d\mu}.$ The space $\mathcal{C}(\mathcal{X})$ denotes continuous real-valued functions on $\mathcal{X}$, and $\mathcal{C}^{k}(\mathcal{X})$ denotes the $k$-times continuously differentiable functions. The Banach space $\mathcal{L}(\mathcal{H}_{1},\mathcal{H}_{2})$ denotes bounded linear operators $\mathsf{T}:\mathcal{H}_{1}\to\mathcal{H}_{2}$, and the operator norm is $\|\mathsf{T}\|_{\mathcal{L}}=\sup_{\|x\|=1}\|\mathsf{T}x\|.$ The adjoint of $\mathsf{T}$ is written $\mathsf{T}^{*}$, and $\mathrm{id}$ denotes the identity operator. An operator is self-adjoint when $\mathsf{T}=\mathsf{T}^{*}$, and it is positive semidefinite when $\mathsf{T}\succeq 0$. For a self-adjoint operator $\mathsf{T}$, $\lambda_{\min}(\mathsf{T})$ and $\lambda_{\max}(\mathsf{T})$ denote its minimal and maximal eigenvalues. The set $\mathcal{P}(\mathcal{X})$ denotes Borel probability measures on $\mathcal{X}$. For $\rho\in\mathcal{P}(\mathcal{X})$, $\operatorname{supp}(\rho)$ denotes the support and $\delta_{x}$ denotes the Dirac mass at $x$. We write $\mathbb{E}[\cdot]$ for expectation and $\mathbb{E}[\cdot|\cdot]$ for conditional expectation. We write $a_{n}=\mathcal{O}(b_{n})$ when $\sup_{n}|a_{n}/b_{n}|<\infty$.
@@ II-B Dynamical System and Flows
In SFL, the flow is the primary regression object: we view the problem of learning vector-fields from trajectory data as identifying the one-step map $(x,h)\mapsto\Phi^{h}(x)$ in an $L^{2}$ space and then lift the flow estimator to the vector field. This choice makes the feature-label pairs well-posed under irregular sampling and enables us to apply spectral regularization methods in learning theory to system identification problems. With this motivation, we now define the dynamical model and its flows. Let $\mathcal{X}\subset\mathbb{R}^{n}$ be compact and $\mathcal{B}=\mathcal{B}(\mathcal{X})$ its Borel $\sigma$-algebra. Consider the control‑affine dynamics:
@@
where $\tau^{s}u$ is the time-shift operator of the input signal $u$ by $s$, such that $(\tau^{s}u)(t)\coloneqq u(t+s),\forall t\geq 0$. We write $\Phi_{u}^{t}\coloneqq\Phi_{u}^{t,0}$. Special case: $u\equiv 0\Rightarrow\Phi_{0}^{t}=\phi^{t}$. We denote flows generated by an autonomous vector field $f$ and control-affine dynamics $(f,g)$ as $\phi^{h}(\ \cdot\ ;f)$ and $\Phi^{h}_{u}(\ \cdot\ ;f,g)$, respectively. Furthermore, we use $\Phi^{h}(\cdot)$ when the discussion is agnostic with respect to the existence of a system input $u$, and $\Phi^{h}_{\rho}(\cdot)$ when the input is determined by a design law $\rho$.
@@ II-C Design Law (State Visitation Distribution)
The design law $\rho_{\mathcal{X}}^{\mathrm{design}}\in\mathcal{P}(\mathcal{X})$, specifies where the experiment spends time in $\mathcal{X}$, and $\rho_{\mathcal{X}}^{\mathrm{design}}$ is a Borel probability measure on $(\mathcal{X},\mathcal{B})$. There are multiple ways to construct a design law; we are interested in the continuous-time time-average law. Along a single path $x(t)=\Phi_{u}^{t}(x_{0})$:
@@
where $\delta_{x}$ is the Dirac measure: $\delta_{x}(A)=\mathbf{1}_{A}(x)$, $A\in\mathcal{B}$. The design law $\rho_{\mathcal{X}}^{\mathrm{design}}$ determines coverage of states; identifiability over the identification domain $\mathcal{X}_{\rm ID}$ requires $\mathcal{X}_{\rm ID}\subseteq\operatorname{supp}(\rho_{\mathcal{X}}^{\mathrm{design}})$.
@@
The design law induces the geometry in which we measure errors and therefore governs identifiability. In particular, we can view identifiability as an input-design problem: choose an input policy $u(\cdot)$ so that the induced $\rho_{\mathcal{X}}^{\mathrm{design}}(u)$ places mass on the domain of interest. Moreover, the population risks $\mathcal{E}(\Phi_{\rho})$ and $\mathcal{E}(f_{\rho})$ used later in the problem formulation are squared $L^{2}$-errors with respect to the design weights $\rho^{\rm design}_{\mathcal{X}}\times\rho_{\mathbf{h}}$ and $\rho^{\rm design}_{\mathcal{X}}$, respectively. The requirement of $\mathcal{X}_{\rm ID}\subseteq\operatorname{supp}(\rho_{\mathcal{X}}^{\mathrm{design}})$ ensures that discrepancies on the identification domain are penalized and thus identifiable. When a specific identification algorithm uses inputs explicitly, we overload $x$ as $x\coloneqq(x,u)$ and regard $\rho_{\mathcal{X}}^{\mathrm{design}}$ as the joint law of sampled states and inputs.
@@ II-D Step Law (Time‑Advance Distribution)
Let $\boldsymbol{h}\coloneqq(0,h_{\max}]$; the step law $\rho_{\boldsymbol{h}}\in\mathcal{P}\big(\boldsymbol{h}\big)$ specifies how far we advance in time when forming one‑step labels $(x_{k},h_{k})\mapsto x_{k+1}$. It is a probability measure on the interval $(0,h_{\max}]$. A step law $\rho_{\boldsymbol{h}}$ can be constructed through index‑average:
$\lim_{N\to\infty}\bigg\langle\varphi,\frac{1}{N}\sum_{k=0}^{N-1}\delta_{h_{k}}\bigg\rangle=\langle\varphi,\rho_{\boldsymbol{h}}\rangle,\quad\forall\ \varphi\in\mathcal{C}(\boldsymbol{h}).$
When modeling a step-window of length $M$, we consider a step-window $h_{1},\dots,h_{M-1}]$ together with the window anchor, $h_{\rm anc}$. Let the bounded step-ratio vector be: $\zeta(M-1)\coloneqq[\zeta_{1},\dots,\zeta_{M-1}]\in\boldsymbol{\zeta}\subset(0,\bar{\zeta}]^{M-1},$ where $\zeta_{k}\coloneqq h_{k}/h_{k-1}$ and $\zeta_{M-1}=h_{M}/h_{M-1}$. We define a $M$-steps window as a tuple $T_{M}\coloneqq(h_{\rm anc},\zeta(M-1))\in\boldsymbol{h}\times\boldsymbol{\zeta}\eqqcolon\mathcal{T}$, with $\rho_{\mathcal{T}}\in\mathcal{P}(\mathcal{T})$. We are ready to formulate the problem of identifying the flow and vector field FS-HP bounds in the SFL framework.
@@ II-E Problem Formulation (FS-HP bound)
$f_{\rho}(x)\coloneqq\lim_{\epsilon\downarrow 0}\mathbb{E}\left[\frac{Y-X}{H}\bigg|X=x,0<H\leq\epsilon\right],$
Under the flow source condition $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$, [11, Eq. 9], and a spectral filter ${\mathsf{g}}_{\lambda}$ with qualification $\nu\geq r$ and Lipschitz exponent $\mu$ (set $\beta\coloneqq\max\{1,2\mu\}$[11, Def. 1], our goal is to establish finite-sample, high-probability bounds of the form:
@@
$\underbrace{\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}}_{\text{vector-field excess risk}}$ $\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}(h)}\log\frac{4}{\eta}\ell^{-\frac{2r}{2r+\beta}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}(h)}\mathbb{E}[H^{2p+2}],$ (2)
which holds with probability at least $1-\eta$ over the model $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}$. Fig. 1 shows the road-map to prove the FS-HP bounds for the flow excess risk in Lemma 3 and the FS-HP bounds for the vector field excess risk in Theorem 1.
vLMM Assumptions Spectral Regularization Assum. III.1 bounded step ratios/coeffs ($\zeta,\alpha,\beta$) Filter qual./Lipschitz (Def. [11, Def. 1]) Assum. III.2 order-$p$ consistency Source cond. on vector field $\Phi_{\rho}\in\Omega^{(\mathcal{V})}_{r,R}$ Assum. III.3 ODE regularity Source cond. on flow $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ Lemma 2 Uniform LTE Lemma 1 Observability via $\mathsf{B}$ Thm. 1 Vector-field FS–HP bound Lemma 3 Flow FS–HP bound
Figure 1: Dependency graph linking assumptions, lemmas, and theorems used in the SFL analysis.
@@ III Variable Step Linear Multistep Method (vLMM) Theory
This section records some vLMM facts we use to turn flow-space error into vector-field error within SFL. The role of vLMM in the SFL framework is twofold: (i) it supplies the windowed surrogate that links vector fields to flows through the multistep forcing/aggregation operators (enabling observability), and (ii) it controls the deterministic discretization bias via order-$p$ truncation guarantees. Since, for variable-step methods, consistency depends only on step-ratio and not on absolute time shifts, a step-window is often parameterized by the step-ratio vector $\zeta$ and an anchor step $h_{\mathrm{anc}}$ in the vLMM framework.